OpenStudy (anonymous):

n inverted right circular conical tank has an altitude equal to 7 inches and a base with radius 1.25 inches. Find the dimensions and the volume of a similar tank whose volume is three times as much as the first conical tank.

OpenStudy (anonymous):

Volume of the conical tank = 13πr2h=13π(1.25)2∗7=11.45in3

OpenStudy (anonymous):

Now, the other tank is similar ==> so its dimensions must be proportional to the dimensions of the original tank ==> let the radius of other tank = 1.25K ===> height of the tank = 7K on solving, we get K = 1.442 new dimensions: radius = 1.25K = 1.25*1.442 = 1.803 inches, height = 7K = 7*1.442 = 10.096 inches

OpenStudy (anonymous):

how did you get the K value?

OpenStudy (anonymous):

Oops! Apparently Open Study doesn't support the symbol, one sec!

OpenStudy (anonymous):

just the k value, that i dont understand where you get it :D

OpenStudy (anonymous):

So, volume of this tank = 13π(1.25K)2(7K) = 3*volume of original tank = 3*11.45 = 34.35in3 That goes above "on solving